One number is 2 times another. If the sum of their reciprocals is 1/5, find the two numbers. Define variables, set up equation and solve.
Answer by checkley71(8403)

X & 2X ARE THE NUMBERS. THUS:
1/X+1/2X=1/5
(2+1)/2X=1/5
3/2X=1/5 NOW CROSS MULTIPLY.
2X=15
X=15/2
X=7.5 ANSWER.
1/7.5+1/15=1/5
(2+1)/15=1/5
3/15=1/5
1/5=1/5